OpenStudy (abbycross167):

Could someone please check my questions? My answers are under the questions... and any help will be greatly appreciated 34. Matty jogs 9 km/hr. Identify the correct conversion factor setup required to compute Matty's speed in m/s. a. 9 km x 1000 m x 1 hr x 60 sec 1 hr 1 km 60 min 1 min b. 9 km x 1000 m x 1 hr x 1 m My Answer~ b

OpenStudy (abbycross167):

I apologize, but not all the answers came up with the post... here are the rest of the answers. c. 9 km x 1000 m x 1 hr x 60 sec 1 hr 9 km 60 min 1 min d. 9 km x 1000 m x 1 hr x 1 min 1 hr 1 km 60 min 60 sec

OpenStudy (anonymous):

The way it's written out is a bit confusing to understand so I'll re-write what the correct way should look like and you can check if you got the same answer

OpenStudy (anonymous):

\[\frac{9 km }{ h }\times \frac{ 1hr }{ 60\min }\times \frac{ 1\min }{ 60\sec }\times \frac{ 1000m }{ 1km }=?\]

OpenStudy (anonymous):

So that the units cancel out and you're left with m/s

OpenStudy (abbycross167):

oh ok, I'm unsure of what the answer would be then :/

OpenStudy (anonymous):

Are your units written in fraction form?

OpenStudy (anonymous):

like min/s or hr/min...etc

OpenStudy (triciaal):